OpenStudy (anonymous):

Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (–3, 0) and co-vertex at (0, 2)

OpenStudy (anonymous):

@dovakiin

OpenStudy (anonymous):

I'm actually doing a whole assigment so yeah :( need some help..

OpenStudy (anonymous):

you have a =+-3, b= =-2 just apply the formula \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (anonymous):

what 's else?

OpenStudy (anonymous):

Oh okay. What are the foci of the ellipse? Graph the ellipse.

OpenStudy (anonymous):

x^2/49+y^2/64=1

OpenStudy (anonymous):

you have a formula for this, too. foci c ^2 = a^2-b^2

OpenStudy (anonymous):

hey, a= 3--> a^2 =9, not 49

OpenStudy (anonymous):

b=2 -->b^2 =4 only

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

c^2 = 9 -4 =5 ---> c = =-sqr (5) no way to have the integer. just estimate between 2 and 3 on the x-axis

OpenStudy (anonymous):

gotcha ok i understand noww

OpenStudy (anonymous):